### v2

@@ 5.3 Domain generalization
Pre-trained vision-language models exhibit a robust ability for domain generalization ability CoOP . We assess the robustness of the proposed RePrompt model on out-of-distribution (OOD) datasets.
@@
Datasets. We follow CoOp CoOP and employ five datasets, namely, ImageNet, ImageNet V2 imagenetv2_paper , ImageNet-Sketch imagenet-sketch_paper , ImageNet-A imagenet-a_paper and ImageNet-R imagenet-r_paper ), to evaluate the generalization ability of RePrompt for out-of-distribution (OOD) data. Following the protocol, we train the model on the ImageNet dataset (as the source dataset) using a 16-shot setting and evaluate its performance on other domain-shifted datasets (as the target datasets). Consequently, we utilize the retrieval database from the ImageNet 16-shot experiment as the retrieval database for the target datasets.
@@
Results. Table 1 summarizes the OOD experiment results, including the accuracy on both source datasets and target datasets. RePrompt achieves optimal results on ImageNet V2 and ImageNet-Sketch while demonstrating performance comparable to PromptSRC promptsrc on ImageNet-R. These results indicate that Reprompt has better generalization for datasets with domain shifts.
Methods ImagNet V2imagenetv2_paper imagenet-sketch_paper imagenet-a_paper Rimagenet-r_paper
CoOP CoOP 71.51 64.20 47.99 49.71 75.21
CoCoOP CoCoOP 71.02 64.07 48.75 50.63 76.18
UPT UPT 72.63 64.35 48.66 50.66 76.24
Maple MaPLe 70.72 64.07 49.15 50.90 76.98
PromptSRC promptsrc 71.27 64.07 49.55 50.90 77.80
RePrompt 74.53 66.66 49.56 49.77 77.48
[…]
Model MVImgNet
$1$-shot $2$-shot $4$-shot $8$-shot $16$-shot Ave.
Vanilla CLIP CLIP 51.4 51.4 51.4 51.4 51.4 51.4
CoOP*  CoOP 62.5 67.8 75.3 79.8 84.0 73.9
VLPT 65.2 69.7 76.8 80.3 84.9 75.3
RePrompt 65.9 73.6 81.1 85.7 90.8 79.4
Table 2: Main results on MVImgNet under the few-shot multi-view settings.RePrompt achieved significant performance improvements due to its retrieval-enhanced module.
@@
CoOPCoOP 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
CoCoOPCoCoOP 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
MapleMaPLe 70.72 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69 66.30
PromptSRCpromptsrc 71.27 93.60 90.25 65.70 70.25 86.15 23.90 67.10 46.87 45.50 68.75 65.81
@@
Table 3: Comparison of RePrompt with existing advanced approaches on cross-dataset benchmark evaluation. The source model (ViT-B/16) is trained on ImageNet imagenet_paper . RePrompt achieves the best results on 5 of 10 datasets. The underline represents the second best.
@@
Vanilla CLIP CLIP 41.9 41.9 41.9 41.9 41.9 63.6 63.6 63.6 63.6 63.6 2.7 2.7 2.7 2.7 2.7
ActionCLIP wang2021actionclip 54.3 56.2 59.3 66.1 59.0 76.7 80.4 87.6 91.8 84.1 4.8 6.9 9.1 12.3 8.3
XCLIP ni2022expanding 60.5 66.8 69.3 71.7 67.1 89.0 91.4 94.7 96.3 92.9 6.6 7.8 9.9 13.7 9.5
A5 ju2022prompting 46.7 50.4 61.3 65.8 56.1 76.3 84.4 90.7 93.0 86.1 4.5 6.7 7.2 9.5 7.0
ViFi-CLIP* rasheed2023fine 61.3 65.3 68.1 70.1 66.4 90.4 92.9 94.4 95.7 93.3 7.2 8.1 10.2 13.9 9.7
@@
Table 4: We compare RePrompt with approaches that explicitly adapt CLIP. ViFi-CLIP only adds an average pooling layer on top of the final features to obtain video-level embeddings. The underline represents the second best.
5.4 Other few-shot classification
@@
Datasets. MVImgNet MVImgNet is a 3D generic dataset that comprises multi-view images, which is a soft bridge between 2D and 3D. It contains 6.5 million frames from 219188 videos, accompanied by comprehensive annotations. These frames covers real-life objects across from 238 classes. Nevertheless, MVImgNet displays a long-tailed distribution of data. Consequently, we eliminate categories with fewer than 16 samples and samples with fewer than 28 frames. In the end, we obtained a subset of MVImgNet, consisting of a total of 220 classes. We adopt the same few-shot setting as used in few-shot image classification. The video understanding task covers HMDB-51 kuehne2011hmdb , UCF-101 ucf101_paper and SSv2 goyal2017ssv2 , where each dataset consists of 2/4/8/16-shot training sets with full categories. The model are evaluated on the first validation split for HMDB-51, UCF-101, and the whole validation split for SSV2.
Training details. We adopt a sparse frame-sampling strategywang2016tsn and configure the number of frames as 16 for multi-view classification and 32 for video understanding. For the video understanding task, models are pre-trained on Kinetics-400k400 to bridge the modality gap and follow a single-view inference rasheed2023fine . he retrieval setup aligns with that of few-shot image classification. However, the training hyperparameters remain consistent with those of ViFi-CLIP-Prompting rasheed2023fine . The minimum retrieval unit in the database is a frame-level embedding, aligned with the input data. The weights of the “Rb adapter” are sourced from another database, where the minimum retrieval unit is a video-level embedding. The use of pooling for simple frame-level temporal aggregation enables the exchange of temporal and multi-view cues in CLIP representations.
Results. Table 2 demonstrates the RePrompt consistently improves performance as the number of shots increases in multi-view classification. For example, it achieves a gain of $+4.2\%$ compared to VLPT. This improvement can be attributed to the retrieval process, which brings more reference information from different perspectives. Table 4 illustrates the effectiveness of RePrompt compared to other CLIP-based approaches for videos. Notably, RePrompt achieves larger gains in low-shot data scenarios, where models are more susceptible to overfitting.
@@
5.5 Ablation
@@
Maple*MaPLe 86.20 95.20 93.57 87.47 37.47 67.53 93.77 74.33 81.20 74.67 71.97 78.49
Tip-Adaptertip-adapter 78.09 95.09 94.62 86.52 39.68 65.90 91.98 75.50 77.93 72.09 70.75 77.10
Tip-Adapter-F*tip-adapter 88.74 95.94 96.79 87.28 46.32 72.87 92.89 84.04 84.38 76.53 73.32 81.74
@@
+Rb adapter 92.91 96.51 97.16 87.42 50.32 73.70 93.76 85.04 86.47 77.54 74.53 83.21
CaFO-F*zhang2023prompt 91.72 97.28 97.97 87.15 54.82 71.34 93.89 83.85 86.10 76.62 74.48 83.20
Reprompt+synt. data 91.95 96.80 98.29 87.23 55.21 76.18 93.92 87.55 87.34 77.39 74.23 84.19
Table 5: Component ablation studies over 11 datasets with 16-shot setting. The average accuracy of RePrompt has been steadily improved through the gradual introduction of retrieval enhancement modules.
Figure 7: Visualization of different learned visual feature manifolds via t-SNE. From left to right, we have CLIP, VLPT, and RePrompt. The image features extracted by RePrompt are more discriminating along the z-axis.
Component ablation. We investigate the effectiveness of RePrompt and report results in Table 5. A progressive approach of incorporating retrieval modules is employed to demonstrate the efficacy of enhancement measures at various stages. The average accuracy of RePrompt consistently improves as we incrementally introduce retrieval augmentation: Retrieval-guiding training denoted as $+$Rg training loss, Retrieval-enhanced visual prompt denoted as $+$Re visual prompt, and Retrieval-based adapter denoted as $+$Rb adapter. We also evaluated the Tip-adapter, which solely utilizes the retrieved samples as prototypes. The findings indicate that retrieval-enhanced visual prompts significantly enhance retrieval-based classification results.
@@
To further substantiate these observations, we employ t-SNE tSNE to visualize the 3D manifold of test features extracted from the CLIP, VLPT, and RePrompt models on the EuroSAT dataset. The t-SNE visualization results, presented in Figure 7, clearly depict that in the high-dimensional classification space, RePrompt achieves a much more distinct separation of image features, with feature distributions more aligned along the z-axis, distinguishing between various categories. Conversely, Tip-Adapter-F, constrained by a frozen visual encoder, exhibits t-SNE visualization results that closely resemble those of CLIP, with considerable overlap in feature distribution. This overlap elucidates why Tip-Adapter performs suboptimally on remote sensing data such as EuroSAT.
@@
Discussions on ReConv and retrieved samples. We explore the impact of the network structure, which is responsible for integrating tokens and queries, on the performance outcomes of the ReConv model. To this end, we conduct a supplementary experiment on ImageNet, where we substitute the convolution layer in ReConv with different layers. ReRNN is a variant in which the convolution layer is replaced with an LSTM layer LSTM ; RNN . ReMLP replaces the convolution layer with a multi-layer perceptron layer. Additionally, the quality of high-retrieval samples also contributes to performance; hence, we examine the effect of utilizing original features and random sample features as inputs. ReConv($z_{q}$) is a variant that exclusively uses $z_{q}$ as input, whereas ReConv+Random employs random sample features. The experimental setup for these variants aligns with that of ReConv, maintaining a comparable number of parameters across both models.
As demonstrated in Table 6, ReConv marginally outperforms the other variants across all shot levels. These results suggest that the specific network structure used for token and query fusion may not significantly affect the overall model performance. Instead, the choice of network structure should take into account the computational cost and optimization challenges, particularly as task complexity increases. However, the quality of retrieval samples plays a crucial role, implying that expanding the search library and enhancing the search methodologies are of greater importance.
@@
Table 6: Ablation study of the visual prompt learner across different input strategies and network structures. The first two rows showcase results using RNN and MLP structures, respectively. The middle two rows illustrate the impact of using original image features and random sample features as inputs. ReConv consistently outperforms the other variations, demonstrating its effectiveness across varying shot scenarios.
@@
[…]
CoOP CoOP 71.20 0.4 14h40min 200
Tip-Adapter-F tip-adapter 73.32 8.19 5min 20
@@
Table 7: Comparison of training efficiency for different methods on 16-shot ImageNet. All experiments are trained with batch 16 on one RTX309 GPU.
@@
CoOP CoOP 71.20 299.64 162.5
Tip-Adapter-F tip-adapter 73.32 10.5 42.5
@@
Table 8: Comparison of inference efficiency for different methods on 16-shot ImageNet. All experiments are trained with batch 16 and tested with batch size 32 on one RTX309 GPU.
6 Retrieval Discussions
Given that the retrieval database is constructed from training data in a few-shot setting, the performance enhancements observed with RePrompt are intimately linked to data statistics. To delve deeper, we investigate the correlation between the efficacy of retrieval enhancement and the intra-class variance of visual features. This specific relationship is depicted in Figure 8, illustrating the association between intra-class visual variance and performance improvement. Aside from the outliers—“StanfordCars” and “DTD”—the remaining datasets exhibit a discernible pattern. Notably, the labels for StanfordCars correspond to vehicle codes, and those for DTD to texture descriptions, categories where CLIP’s classification capabilities manifest notable deficiencies. Consequently, results from these two datasets are excluded from the general analysis to delineate retrieval enhancement patterns more accurately.
@@ 6 Retrieval Discussions
6.1 Analysis of retrieval
Experiments are conducted across FGVC Aircraft, Oxford Pets and ImageNet-1K to quantitatively assess the influence of retrieval in various few-shot scenarios. FGVC Aircraft has the smallest intra-class visual variance, while Oxford Pets has the largest; ImageNet-1K falls intermediate between these two. By comparing the optimal retrieval parameter adjustments among these datasets, we can analysis the influence of intra-class visual variance on the retrieval mechanism.
The increasing value of the retrieval-enhanced prompt number $k_{\mathrm{re}}+2$ indicates that RePrompt benefits from additional retrieved information. As illustrated in Table 9, the optimal $k_{\mathrm{re}}$ parameter varies among the three datasets. Specifically, we observe that $k_{\mathrm{re}}=14$ consistently underperforms across all few-shot settings for Oxford Pets. This suggests that the retrieved information is detrimental and lacks confidence in contexts with significant intra-class visual variance. In contrast, $k_{\mathrm{re}}=14$ demonstrates improved average performance on FGVC Aircraft, underscoring its utility in scenarios with lesser intra-class variance.
@@ 6.1 Analysis of retrieval
Table 9: Ablation study on different visual prompt numbers over Fgvc Aircraft and Oxford Pets with few-shot settings. The retrieval-enhanced prompt is mainly affected by the convolution kernel size of REConv.
The factor $n$ determines the number of retrieval samples, $\left|\boldsymbol{K}\right|=C\times\boldsymbol{n}$, used in retrieval-guiding training loss. Increasing the value of $n$ signifies a more relaxed constraint on retrieval. As shown in Table 10, $\boldsymbol{n}=1$ achieves optimal performance in the 16-shot setting on Oxford Pets, which implies that RePrompt prefers the output of CLIP with retrieval-enhanced prompt. However, in low-shot scenarios(e.g., $\boldsymbol{n}=2/1$ under 2/1-shot setting ), RePrompt may require more additional references due to the scarcity of training data.
@@
Table 10: Ablation studies on retrieving $n$ samples for each class over Fgvc Aircraft and Oxford Pets with few-shot settings. The $k$NN-guiding training is weak in low-shot settings, since the model may require reference on the $k$-nn classifier.
Figure 9: Tuning retrieval parameter results over 3 datasets under the 16-shot learning setting. From top to bottom, images report ablation experiments about tuning retrieval parameters on FGVC Aircraft, ImageNet, and Oxford Pets, respectively.

### v3

@@ V-B Domain Generalization
Pre-trained vision-language models exhibit a robust ability for domain generalization ability [6]. We assess the robustness of the proposed RePrompt model on out-of-distribution (OOD) datasets.
@@
Datasets. We follow CoOp [6] and employ five datasets, namely, ImageNet, ImageNet V2 [60], ImageNet-Sketch [61], ImageNet-A [62] and ImageNet-R [63]), to evaluate the generalization ability of RePrompt for out-of-distribution (OOD) data. Following the protocol, we train the model on the ImageNet dataset (as the source dataset) using a 16-shot setting and evaluate its performance on other domain-shifted datasets (as the target datasets). Consequently, we utilize the retrieval database from the ImageNet 16-shot experiment as the retrieval database for the target datasets.
@@
Results. Table II summarizes the OOD experiment results, including the accuracy on both source datasets and target datasets. RePrompt achieves optimal results on ImageNet V2 and ImageNet-Sketch while demonstrating performance comparable to PromptSRC [15] on ImageNet-R. These results indicate that Reprompt has better generalization for datasets with domain shifts. RePrompt struggles on ImageNet-A and ImageNet-R. This is likely due to a lack of diversity in the retrieval database, only ImageNet, which limits generalization to adversarial and artistic images. In practical scenarios, we can improve the deficiency by expanding the search base and improving the search quality.
TABLE I: Comparison of RePrompt with existing advanced approaches on cross-dataset benchmark evaluation. The source model (ViT-B/16) is trained on ImageNet[46]. RePrompt achieves the best results on 5 of 10 datasets. The underline represents the second best.
@@
CoOP[6] 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
CoCoOP[37] 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
Maple[13] 70.72 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69 66.30
PromptSRC[15] 71.27 93.60 90.25 65.70 70.25 86.15 23.90 67.10 46.87 45.50 68.75 65.81
@@
[…]
Methods ImagNet V2[60] S[61] A[62] R[63]
CoOP[6] 71.51 64.20 47.99 49.71 75.21
CoCoOP[37] 71.02 64.07 48.75 50.63 76.18
UPT[12] 72.63 64.35 48.66 50.66 76.24
Maple[13] 70.72 64.07 49.15 50.90 76.98
PromptSRC[15] 71.27 64.07 49.55 50.90 77.80
RePrompt 74.53 66.66 49.56 49.77 77.48
TABLE III: The main results on MVImgNet are under the few-shot multi-view settings.RePrompt achieved significant performance improvements due to its retrieval-enhanced module. The methods marked with “*” are reproduced by us.
Model MVImgNet
$1$-shot $2$-shot $4$-shot $8$-shot $16$-shot Ave.
Vanilla CLIP[4] 51.4 51.4 51.4 51.4 51.4 51.4
CoOP* [6] 62.5 67.8 75.3 79.8 84.0 73.9
VLPT 65.2 69.7 76.8 80.3 84.9 75.3
RePrompt 65.9 73.6 81.1 85.7 90.8 79.4
TABLE IV: We compare RePrompt with approaches that explicitly adapt CLIP. ViFi-CLIP only adds an average pooling layer on top of the final features to obtain video-level embeddings. The underline represents the second best.
@@
Vanilla CLIP[4] 41.9 41.9 41.9 41.9 41.9 63.6 63.6 63.6 63.6 63.6 2.7 2.7 2.7 2.7 2.7
ActionCLIP[64] 54.3 56.2 59.3 66.1 59.0 76.7 80.4 87.6 91.8 84.1 4.8 6.9 9.1 12.3 8.3
XCLIP[65] 60.5 66.8 69.3 71.7 67.1 89.0 91.4 94.7 96.3 92.9 6.6 7.8 9.9 13.7 9.5
A5[66] 46.7 50.4 61.3 65.8 56.1 76.3 84.4 90.7 93.0 86.1 4.5 6.7 7.2 9.5 7.0
ViFi-CLIP*[67] 61.3 65.3 68.1 70.1 66.4 90.4 92.9 94.4 95.7 93.3 7.2 8.1 10.2 13.9 9.7
@@
V-C Additional Few-shot Classification
@@
Datasets. MVImgNet [68] is a 3D generic dataset that comprises multi-view images, which is a soft bridge between 2D and 3D. It contains 6.5 million frames from 219188 videos, accompanied by comprehensive annotations. These frames covers real-life objects across from 238 classes. Nevertheless, MVImgNet displays a long-tailed distribution of data. Consequently, we eliminate categories with fewer than 16 samples and samples with fewer than 28 frames. In the end, we obtained a subset of MVImgNet, consisting of a total of 220 classes. We adopt the same few-shot setting as used in few-shot image classification. The video understanding task covers HMDB-51 [69], UCF-101 [59] and SSv2 [70], where each dataset consists of 2/4/8/16-shot training sets with full categories. The model are evaluated on the first validation split for HMDB-51, UCF-101, and the whole validation split for SSV2.
Training details. We adopt a sparse frame-sampling strategy [71] and configure the number of frames as 16 for multi-view classification and 32 for video understanding. For the video understanding task, models are pre-trained on Kinetics-400 [72] to bridge the modality gap and follow a single-view inference [67]. he retrieval setup aligns with that of few-shot image classification. However, the training hyperparameters remain consistent with those of ViFi-CLIP-Prompting [67]. The minimum retrieval unit in the database is a frame-level embedding, aligned with the input data. The weights of the “Rb adapter” are sourced from another database, where the minimum retrieval unit is a video-level embedding. The use of pooling for simple frame-level temporal aggregation enables the exchange of temporal and multi-view cues in CLIP representations.
Results. Table III demonstrates the RePrompt consistently improves performance as the number of shots increases in multi-view classification. For example, it achieves a gain of $+4.2\%$ compared to VLPT. This improvement can be attributed to the retrieval process, which brings more reference information from different perspectives. Table IV illustrates the effectiveness of RePrompt compared to other CLIP-based approaches for videos. Notably, RePrompt achieves larger gains in low-shot data scenarios, where models are more susceptible to overfitting.
@@
V-D Ablation Study
TABLE V: Component ablation studies over 11 datasets with 16-shot setting. The average accuracy of RePrompt has been steadily improved through the gradual introduction of retrieval enhancement modules. We reproduce methods marked with “*” on current datasets. “Reprompt+Synthetic data” means introducing synthetic data on the basis of RePrompt.
@@ V-D Ablation Study
Maple*[13] 86.20 95.20 93.57 87.47 37.47 67.53 93.77 74.33 81.20 74.67 71.97 78.49
Tip-Adapter[21] 78.09 95.09 94.62 86.52 39.68 65.90 91.98 75.50 77.93 72.09 70.75 77.10
Tip-Adapter-F*[21] 88.74 95.94 96.79 87.28 46.32 72.87 92.89 84.04 84.38 76.53 73.32 81.74
@@
+Rb adapter(RePrompt) 92.91 96.51 97.16 87.42 50.32 73.70 93.76 85.04 86.47 77.54 74.53 83.21
CaFO-F*[30] 91.72 97.28 97.97 87.15 54.82 71.34 93.89 83.85 86.10 76.62 74.48 83.20
Reprompt+Synthetic data 91.95 96.80 98.29 87.23 55.21 76.18 93.92 87.55 87.34 77.39 74.23 84.19
Figure 6: Visualization of different learned visual feature manifolds via t-SNE. From left to right, we have CLIP, VLPT, and RePrompt. The image features extracted by RePrompt discriminate more along the z-axis.
Component ablation. We investigate the effectiveness of RePrompt and report results in Table V. A progressive approach of incorporating retrieval modules is employed to demonstrate the efficacy of enhancement measures at various stages. The average accuracy of RePrompt consistently improves as we incrementally introduce retrieval augmentation: Retrieval-guided training denoted as $+$Rg training loss, Retrieval-enhanced visual prompt denoted as $+$Re visual prompt, and Retrieval-based adapter denoted as $+$Rb adapter. We also evaluated the Tip-adapter, which solely utilizes the retrieved samples as prototypes. The findings indicate that retrieval-enhanced visual prompts significantly enhance retrieval-based classification.
@@
To further substantiate these observations, we employ t-SNE [73] to visualize the 3D manifold of test features extracted from the CLIP, VLPT, and RePrompt models on the EuroSAT dataset. The t-SNE visualization results, presented in Fig. 6, clearly depict that in the high-dimensional classification space, RePrompt achieves a much more distinct separation of image features, with feature distributions more aligned along the z-axis, distinguishing between various categories. Conversely, Tip-Adapter-F, constrained by a frozen visual encoder, exhibits t-SNE visualization results that closely resemble those of CLIP, with considerable overlap in feature distribution. This overlap elucidates why Tip-Adapter performs suboptimally on remote sensing data such as EuroSAT.
@@
Discussions on ReConv and retrieved samples. We explore the impact of the network structure, which is responsible for integrating tokens and queries, on the performance outcomes of the ReConv model. To this end, we conduct a supplementary experiment on ImageNet, where we substitute the convolution layer in ReConv with different layers. ReRNN is a variant in which the convolution layer is replaced with an LSTM layer [74, 75]. ReMLP replaces the convolution layer with a multi-layer perceptron layer. Additionally, the quality of high-retrieval samples also contributes to performance; hence, we examine the effect of utilizing original features and random sample features as inputs. ReConv($z_{q}$) is a variant that exclusively uses $z_{q}$ as input. The ReConv+Random employs random sample features from the retrieval database. The experimental setup for these variants aligns with that of ReConv, maintaining a comparable number of parameters across both models.
As demonstrated in Table VI, ReConv marginally outperforms the other variants across all shot levels. These results suggest that the specific network structure used for token and query fusion may not significantly affect the overall model performance. Instead, the choice of network structure should take into account the computational cost and optimization challenges, particularly as task complexity increases. However, the quality of retrieval samples plays a crucial role, implying that expanding the search library and enhancing the search methodologies are of greater importance.
TABLE VI: Ablation study of the visual prompt learner across different input strategies and network structures. The first two rows showcase results using RNN and MLP structures, respectively. The middle two rows illustrate the impact of using original image features and random sample features as inputs. ReConv consistently outperforms the other variations, demonstrating its effectiveness across varying shot scenarios.
@@
TABLE VII: Comparison of training efficiency for different methods on 16-shot ImageNet. All experiments are trained with batch 16 on one RTX309 GPU.
@@
CoOP[6] 71.20 0.4 14h40min 200
Tip-Adapter-F[21] 73.32 8.19 5min 20
@@
TABLE VIII: Comparison of inference efficiency for different methods on 16-shot ImageNet. All experiments are trained with batch 16 and tested with batch size 32 on one RTX309 GPU.
@@
CoOP*[6] 71.20 299.64 162.5
Tip-Adapter-F*[21] 73.32 10.5 42.5
@@
[…]
VI Retrieval Discussions
Given that the retrieval database is constructed from training data in a few-shot setting, the performance enhancements observed with RePrompt are intimately linked to data statistics. To delve deeper, we investigate the correlation between the efficacy of retrieval enhancement and the intra-class variance of visual features. This specific relationship is depicted in Fig. 7, illustrating the association between intra-class visual variance and performance improvement. Aside from the outliers—“StanfordCars” and “DTD”—the remaining datasets exhibit a discernible pattern. Notably, the labels for StanfordCars correspond to vehicle codes, and those for DTD to texture descriptions, categories where CLIP’s classification capabilities manifest notable deficiencies. Consequently, results from these two datasets are excluded from the general analysis to delineate retrieval enhancement patterns more accurately.
@@ VI Retrieval Discussions
TABLE IX: Ablation study on different visual prompt numbers over Fgvc Aircraft and Oxford Pets with few-shot settings. The retrieval-enhanced prompt is mainly affected by the convolution kernel size of REConv.
@@
TABLE X: Ablation studies on retrieving $n$ samples for each class over Fgvc Aircraft and Oxford Pets with few-shot settings. The retrieval-guided training is weak in low-shot settings, since the model may require reference on the $k$-nn classifier.
@@
VI-A Analysis of Retrieval
Experiments are conducted across FGVC Aircraft, Oxford Pets and ImageNet-1K to quantitatively assess the influence of retrieval in various few-shot scenarios. FGVC Aircraft has the smallest intra-class visual variance, while Oxford Pets has the largest; ImageNet-1K falls intermediate between these two. By comparing the optimal retrieval parameter adjustments among these datasets, we can analyze the influence of intra-class visual variance on the retrieval mechanism.
The increasing value of the retrieval-enhanced prompt number $k_{\mathrm{re}}+2$ indicates that RePrompt benefits from additional retrieved information. As illustrated in Table IX, the optimal $k_{\mathrm{re}}$ parameter varies among the three datasets. Specifically, we observe that $k_{\mathrm{re}}=14$ consistently underperforms across all few-shot settings for Oxford Pets. In contrast, $k_{\mathrm{re}}=14$ demonstrates improved average performance on FGVC Aircraft, underscoring its utility in scenarios with lesser intra-class variance. This suggests that the retrieved information is detrimental and lacks confidence in contexts with significant intra-class visual variance.
The factor $n$ determines the number of retrieval samples, $\left|\boldsymbol{K}\right|=C\times\boldsymbol{n}$, used in retrieval-guided training loss. Increasing the value of $n$ signifies a more relaxed constraint on retrieval. As shown in Table X, $\boldsymbol{n}=1$ achieves optimal performance in the 16-shot setting on Oxford Pets, which implies that RePrompt prefers the output of CLIP with retrieval-enhanced prompt. In low-shot scenarios(e.g., $\boldsymbol{n}=2/1$ under 2/1-shot setting ), RePrompt may require more additional references due to the scarcity of training data.